A number consists of two digits whose sum is 10 . If 72 be subtracted from the number , the digits are reversed . Find the number ? Sir can u please help me .
Answer by Paul(988)

Let the 1o-digit be x
Let the one-digit be y
Now: x+y=10
y=10-x (subsitution)
Reversed digits: 10y+x
Equation:
10x+y-72=10y+x
Subsitute for y:
10x+(10-x)-72=10(10-x)+x
9x-62=100-9x
18x=162
x=9
y=10-9
y=1
Hence, the to-digit number is 9, and the one digit number is 1 and the whole number is 91.
Paul.